What is the de Broglie wavelength of a 162 g baseball with a speed of 38.0 m/s ? (in m)

What is the speed of a 162 g baseball with a de Broglie wavelength of 0.230 nm ? (in m/s)
Answer #1

lambda = h/p = h/mv

lambda = 6.63*10^-34/0.162*38 = 1.08*10^-34 m

v = h/m*lambda = 6.63*10^-34/0.162*0.23*10^-9 = 177.94*10^-25 m/s